# SeSE: A Structural Information-Guided Uncertainty Quantification Framework for Hallucination Detection in LLMs

### v2

@@ Abstract
Reliable uncertainty quantification (UQ) is essential for deploying large language models (LLMs) in safety-critical scenarios, as it enables them to abstain from responding when uncertain, thereby avoiding “hallucinating” falsehoods. However, state-of-the-art UQ methods primarily rely on semantic probability distributions or pairwise distances, overlooking latent semantic structural information that could enable more precise uncertainty estimates. This paper presents Semantic Structural Entropy (SeSE), a principled UQ framework that quantifies the inherent semantic uncertainty of LLMs from a structural information perspective for hallucination detection. SeSE operates in a zero-resource manner and is applicable to both open- and closed-source LLMs, making it an “off-the-shelf” solution for new models and tasks. Specifically, to effectively model semantic spaces, we first develop an adaptively sparsified directed semantic graph construction algorithm that captures directional semantic dependencies while automatically pruning unnecessary connections that introduce negative interference. We then exploit latent semantic structural information through hierarchical abstraction: SeSE is defined as the structural entropy of the optimal semantic encoding tree, formalizing intrinsic uncertainty within semantic spaces after optimal compression. A higher SeSE value corresponds to greater uncertainty, indicating that LLMs are highly likely to generate hallucinations. In addition, to enhance fine-grained UQ in long-form generation—where existing methods often rely on heuristic sample-and-count techniques—we extend SeSE to quantify the uncertainty of individual claims by modeling their random semantic interactions, providing theoretically explicable hallucination detection. Extensive experiments across 29 model-dataset combinations show that SeSE significantly outperforms advanced UQ baselines, including strong supervised methods and the recently proposed KLE. These findings highlight the importance of semantic structural information for precise UQ in LLMs, establishing SeSE as a reliable framework for assessing the trustworthiness of LLM-based systems [^1] The code and data is available: https://github.com/SELGroup/SeSE.
I Introduction
Large language models (LLMs) have been widely adopted across various fields, including spatio-temporal data modeling [1], sentiment analysis [2] and recommender systems [3, 4, 5], due to their impressive general intelligence. However, a key limitation of LLMs is hallucinations [6]—the generation of plausible yet incorrect statements, which hinders their deeper adoption in real-world [7]. Despite extensive research having sought to eliminate hallucinations by verifying model outputs against evidence retrieved from external knowledge bases [8, 9, 10], they are often domain-specific and fail to address hallucinations beyond pure fact-checking [7]. A promising direction for avoiding hallucinations is uncertainty quantification (UQ) [11, 12]. As shown in Figure 1, UQ mines uncertainty information and assigns an uncertainty score to the model when answering a given input. A low score suggests LLM responses for a given input are acceptable, whereas a high score should trigger LLMs to abstain from responding and warn users of potential hallucinations.
@@ I Introduction
But the open-ended nature of LLM generation hinders the direct application of traditional UQ methods [13], which treat LLM outputs as conventional autoregressive sequences and consider only their lexical uncertainty. Since response correctness fundamentally depends on semantics in most LLM applications, and distinct token sequences can convey identical meanings, uncertainty in the semantic space is a more essential indicator of response trustworthiness than lexical uncertainty. To address this issue, semantic entropy (SE) [14] and its subsequent extensions [15, 16, 17] quantify uncertainty at the semantic rather than token level, achieving state-of-the-art (SOTA) performance in UQ tasks.
However, current semantic UQ methods exhibit three intrinsic limitations. First, they ignore the structural information embedded within semantic spaces. As illustrated in Figure 2 (2), real-world semantic spaces are often hierarchically organized—substructures recursively contain finer sub-substructures. According to the “Compositional Similarity” principle [18], identifying the hierarchical structure enables more precise UQ, as differences among substructures may help distinguish superficially similar semantic spaces. SE is a binary, one-cut measurement that considers only the uncertainty of the distribution over semantic clusters. Kernel language entropy (KLE) [17], a recent follow-up of SE, computes the von Neumann entropy of semantic graph kernels, but the decomposed semantic substructures are non-hierarchical and any substructure can be compared with any other. This violates the “Compositional Similarity” principle and limits the precision of UQ.
Figure 1: Illustration of LLM UQ for avoiding hallucinating.
Second, existing semantic UQ methods often represent semantic spaces using suboptimal undirected, complete graphs [17, 19]. Undirected graphs capture only symmetric similarity between responses, implicitly assuming that the similarity from A to B equals that from B to A. However, in semantic studies, entailment relations are directionally asymmetric [20]; for example, “apples” entails “green apple”, but not vice versa. Ignoring this directionality discards crucial semantic logic and leads to irreversible information loss. Moreover, complete graphs may contain numerous low-value edges (e.g., edges with near-zero weights) that obfuscate core semantic structures and reduce computational efficiency. Existing sparsification approaches primarily fall into two categories: (i) removing redundant inter-cluster edges through pairwise semantic equivalence assessment [16, 17], which has a worst-case time complexity $O(n^{2})$ and incurs substantial computational cost; and (ii) relying on domain knowledge to manually set pruning thresholds [21, 22, 23], which fails to generalize to task-specific semantics.
Thirdly, granular uncertainty estimation for long-form generations remains challenging. Although existing semantic UQ methods have made notable progress in short-form settings where generations typically consist of one or two sentences [24, 14, 16, 15, 17], real-world LLM applications predominantly involve long-form generations that interleave numerous true or false claims. Recent works [25, 26, 22, 27] have begun to evaluate the factual accuracy of individual claim outputs based on the concept of self-consistency [25]. However, they rely purely on heuristic sample-and-count techniques, which fail to capture more sophisticated semantic relationships between claims and responses.
To address these challenges, we present SeSE, a principled zero-resource UQ framework that quantifies semantic uncertainty from the perspective of structural information theory for LLM hallucination detection. SeSE is self-contained: it relies solely on sampled LLM responses and can thus be applied to both open-source and proprietary LLMs without requiring any additional training or external tools. Our framework offers three key advancements: First, we develop an adaptively sparsified directed semantic graph construction algorithm (AS-DSG) to effectively represent semantic spaces. The directed semantic graph, with edges weighted by directional entailment probabilities, captures both semantic directionality and similarity. Inspired by neighbor network construction in cancer cell analysis [28], AS-DSG employs an adaptive graph sparsification mechanism that minimizes the one-dimensional structural entropy to prune the directed graph into an optimal $k$-nearest neighbor (kNN) graph, thereby reducing redundancy. This procedure does not require manually setting $k$ or costly pairwise comparisons, adaptively prunes trivial edges, and preserves core semantic structure. Then, we fully capture semantic structural information by constructing the hierarchical abstraction using an encoding tree. SeSE is defined as the structural entropy of the $K$-dimensional optimal encoding tree derived from the constructed semantic graph based on the principle of structural entropy minimization [29]. SeSE thus quantifies the residual uncertainty within semantic spaces after optimal compression. Semantic spaces with clear regularities exhibit low entropy (low uncertainty), whereas disordered spaces resist compression and yield high entropy (high uncertainty). Furthermore, we extend SeSE to quantify claim-level uncertainty in long-form generation, improving granular UQ by considering fine-grained semantic dependencies within semantic spaces. Following prior work on long-form generation decomposition [27, 30], we segment LLM outputs into atomic claims and construct a response-claim bipartite graph to model their semantic interactions. The SeSE of a claim is defined as the uncertainty of reaching that claim through random walks on the response-claim graph, offering both theoretical and empirical advantages over conventional sample-and-count techniques. Claims with low SeSE lie within core semantic regions, implying they are frequently accessed during the LLM’s sequence generation and are thus highly likely to be factual. Conversely, claims with high SeSE occupy peripheral semantic positions, increasing their likelihood of being associated with hallucinations. In summary, the main contributions of this paper are:
@@
  • We propose SeSE, a principled and self-contained UQ framework applicable to both open- and closed-source LLMs. To the best of our knowledge, this is the first work to incorporate semantic structural information into UQ.
  • We develop an adaptively sparsified directed semantic graph construction algorithm (AS-DSG) that captures directional semantic relationships while adaptively pruning insignificant edges to reduce interference.
  • Based on structural information theory, we quantify the inherent uncertainty within the semantic space of LLMs. We refine naive structural entropy to enable hierarchical abstraction under asymmetric connections by introducing an optimized redefinition and an associated optimization operator.
  • We extend SeSE to provide enhanced granular UQ in long-form generation by accounting for fine-grained semantic interactions, offering both theoretical and empirical advantages over existing heuristic techniques.
  • We conducted comprehensive experiments across 29 scenarios, encompassing both open- and closed-source LLMs and multiple free-form QA datasets with varying generation lengths. Experimental results highlight SeSE’s superior performance over existing SOTA UQ baselines.
II Preliminaries and Problem Formulation
In this section, we first define the LLM uncertainty quantification problem in Section II-A, then present the basic concepts of predictive uncertainty (Section II-C) and naive structural entropy (Section II-B).
II-A Problem Statement
@@
Given a pre-trained LLM $\mathcal{M}$, SeSE takes as input the user query $x$ and $N$ responses $\mathcal{R}(\cdot|X)=\left\{r_{T=t}^{1},\ldots,r_{T=t}^{N}\right\}$ that are sampled from $\mathcal{M}$ at temperature $T=t$, and outputs an uncertainty score $U(x)$ that is positively correlated with the probability of the best generation (i.e., greedy-decoded response) $r_{T=0}$ being hallucination. LLM-based systems can use $U(x)$ as a quantitative indicator to assess the trustworthiness of $\mathcal{M}$’s response to $x$.
II-B Naive Structural Entropy
Structural entropy is an extension of Shannon entropy for measuring the uncertainty of a graph under hierarchical abstraction. It represents the minimum code length required to describe a graph’s hierarchy during random walks [31]. An encoding tree with the minimum structural entropy captures the natural structure of the semantic space.
Definition 1 (Encoding tree).
@@ II-B Naive Structural Entropy
Given the graph $G$, the encoding tree $\mathcal{T}$ for $G$ is a rooted tree satisfying the following conditions: (1) The root node $\lambda$ contains all nodes in $G$, $\mathcal{T}_{\lambda}=V$. Each node $\alpha\in\mathcal{T}$ represents a partition of responses $\mathcal{T}_{\alpha}\subseteq V$. For any leaf node $\gamma$, $\mathcal{T}_{\gamma}$ contains a single response from $V$. (2) Each non-leaf node $\alpha$ has a nonempty set of immediate successors denoted as $\beta_{0},\beta_{1},\ldots,\beta_{l}$. The collection of subsets $\{\mathcal{T}_{\beta_{0}},\mathcal{T}_{\beta_{1}},\ldots,\mathcal{T}_{\beta_{l}}\}$ forms a partition of $\mathcal{T}_{\alpha}$.
Definition 2 (One-dimensional structural entropy).
The one-dimensional structural entropy of $G$ reflects the dynamical uncertainty of $G$ based on random walks, defined as:
@@
Definition 3 (Structural Entropy of the Encoding Tree $\mathcal{T}$).
@@
$H^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{{vol}(G)}\log_{2}\frac{\mathcal{V}_{\alpha}}{\mathcal{V}_{\alpha^{-}}},$ (3)

### v3

# SeSE: Black-Box Uncertainty Quantification for Large Language Models Based on Structural Information Theory

@@ Abstract
Reliable uncertainty quantification (UQ) is essential for deploying large language models (LLMs) in safety-critical scenarios, as it enables them to abstain from responding when uncertain, thereby avoiding hallucinations, i.e., plausible yet factually incorrect responses. However, while semantic UQ methods have achieved advanced performance, they overlook latent semantic structural information that could enable more precise uncertainty estimates. In this paper, we propose Semantic Structural Entropy (SeSE), a principled black-box UQ framework applicable to both open- and closed-source LLMs. To reveal the intrinsic structure of the semantic space, SeSE constructs its optimal hierarchical abstraction through an encoding tree with minimal structural entropy. The structural entropy of this encoding tree thus quantifies the inherent uncertainty within LLM semantic space after optimal compression. Additionally, unlike existing methods that primarily focus on simple short-form generation, we extent SeSE to provide interpretable, granular uncertainty estimation for long-form outputs. We theoretically prove that SeSE generalizes semantic entropy, the gold standard for UQ in LLMs, and empirically demonstrate its superior performance over strong baselines across 24 model-dataset combinations.
1 Introduction
Large language models (LLMs) have been widely adopted across various fields [Perplexity, 2025, Liu et al., 2025a], owing to their impressive general intelligence. However, even state-of-the-art (SOTA) models frequently generate plausible yet incorrect statements [OpenAI, 2025], a phenomenon known as hallucinations [Huang et al., 2025], which impedes their deployment in risk-sensitive domains. Despite extensive research seeking to eliminate hallucinations [Tian et al., 2023a, Rafailov et al., 2023], complete solutions remain elusive. A promising approach for avoiding hallucinations is uncertainty quantification (UQ) [Lin et al., 2024], which estimates the likelihood of an LLM hallucinating falsehoods for a given input [Cole et al., 2023]. Lower uncertainty suggests the initial response is acceptable, whereas higher uncertainty should trigger LLMs to abstain from answering and alert users to potential errors.
@@ 1 Introduction
However, the open-ended nature of LLM generation hinders the direct application of traditional UQ methods [Liu et al., 2020, Malinin and Gales, 2021], which treat LLM outputs as autoregressive sequences and consider only lexical uncertainty. Since response correctness fundamentally depends on semantics and distinct token sequences can convey identical meanings, uncertainty in the semantic space serves as a more reliable indicator of trustworthiness than lexical uncertainty. To address this issue, Semantic Entropy (SE) [Farquhar et al., 2024] was proposed to quantify uncertainty at the semantic rather than token level, serving as the gold standard for UQ in LLMs [Huang et al., 2025].
Despite its success, SE and its subsequent extensions [Farquhar et al., 2024, Nikitin et al., 2024, Qiu and Miikkulainen, 2024, Li et al., 2025b] suffer from key limitations. First, they fail to capture the inherent semantic structure that defines the organizational principle of the semantic space, which could enable more precise uncertainty estimates Li et al. [2024]. As illustrated in Figure 1 (2), semantic spaces are often hierarchically organized, with substructures recursively containing finer sub-substructures. According to the “Compositional Similarity” principle [Boiman and Irani, 2006], identifying this hierarchy is essential, as differences between substructures could help distinguish superficially similar semantic spaces. Second, while current semantic UQ methods have progressed in short-form settings where generations typically consist of one or two sentences [Farquhar et al., 2024, Nikitin et al., 2024, Qiu and Miikkulainen, 2024, Li et al., 2025b], they often lack sufficient granularity for real-world applications where LLMs often generate long-form paragraphs comprising interwoven true and false claims. Recent works [Manakul et al., 2023, Mohri and Hashimoto, 2024, Zhang et al., 2024, Jiang et al., 2024] have begun to quantify claim-level uncertainty based on the concept of self-consistency [Manakul et al., 2023]. However, they primarily rely on heuristic sample-and-count techniques, which offer limited interpretability and fail to capture fine-grained semantic dependencies between claims and responses.
To address these issues, we propose Semantic Structural Entropy (SeSE), a principled black-box UQ framework that does not require access to LLM internal states and is applicable to both open- and closed-source LLMs. For the first issue, to represent the hierarchical structure of LLM semantic space, SeSE formalizes its hierarchical abstraction by constructing an optimal $K$-dimensional encoding tree adhering to the structural entropy minimization principle [Li and Pan, 2016]. SeSE then quantifies the inherent uncertainty after optimal compression using the structural entropy of this tree. For the second issue, we extend SeSE to provide interpretable, granular claim-level uncertainty estimation in long-form generation. Following prior decomposition methods [Min et al., 2023], we segment long-form outputs into atomic claims and construct a claim-response bipartite graph to capture fine-grained semantic dependencies. The SeSE of a claim is defined as the uncertainty of reaching that claim via random walks on this graph. Importantly, we theoretically prove that SeSE can recover SE where the encoding tree is restricted to a single layer ($K=1$). In summary, the contributions of this paper are:
@@
  • We propose SeSE, a principled black-box UQ framework built upon structural information theory and applicable to both open- and closed-source LLMs. To the best of our knowledge, this is the first work to incorporate semantic structural information into UQ for LLMs.
  • We extend SeSE to provide interpretable, granular claim-level UQ in long-form generation by modeling random semantic interactions within claim-response bipartite graphs.
  • We prove that SeSE is a generalization of semantic entropy.
  • We empirically compare SeSE against advanced baselines across 24 model-dataset pairs, achieving SOTA results.
2 Preliminaries
Problem Formulation
@@ 2 Preliminaries
Given a pre-trained LLM $\mathcal{M}$, SeSE takes as input a query $x$ and $N$ responses $R(\cdot|x)=\left\{r^{1}_{T=t},\ldots,r^{N}_{T=t}\right\}$ generated from $\mathcal{M}$ at temperature $T=t$, and outputs a relative uncertainty score $U(x)$ that is positively correlated with the probability of the model’s most confident (greedy-decoded) response $r_{T=0}$ being hallucinatory. LLM-based systems can use $U(x)$ as a quantitative indicator to assess the trustworthiness of $\mathcal{M}$’s response to $x$ and decide whether to abstain in high-uncertainty cases.
It is important to note that $U(x)$ is not an exact probability of model correctness, the latter pertaining to the orthogonal topic of model calibration Liu et al. [2025b]. Following prior work, we focus on estimating total uncertainty, which comprises epistemic and aleatoric uncertainty Hou et al. [2024], as they jointly contribute to hallucinations.
Semantic Uncertainty
To measure uncertainty at the semantic rather than the token level, Farquhar et al. [2024] introduced Semantic Entropy (SE). Given an input $x$ and the set of all possible semantic clusters $\Omega$, SE is defined as:
$\text{SE}(x)=-\sum_{C\in\Omega}p(C\mid x)\log p(C\mid x).$ (1)
Since $\Omega$ is typically intractable, SE is estimated using $M$ clusters $\{C_{i}\}_{i=1}^{M}$ extracted from generated samples:
$\text{SE}(x)\approx-\sum_{i=1}^{M}p(C_{i}\mid x)\log p(C_{i}\mid x),$ (2)
where $p(C_{i}\mid x)$ is the normalized semantic probability. When token-level likelihoods are unavailable, $p(C_{i}\mid x)$ can be approximated by the frequency of samples falling into each cluster: $p(C_{i}\mid x)\approx\frac{1}{N}\sum_{j=1}^{N}\mathbb{I}(S_{j}\in C_{i})$, which is referred to as Discrete Semantic Entropy (DSE).
Hierarchical Abstraction
Inspired by the concept of the partitioning tree Li and Pan [2016], we construct the hierarchical abstraction (Definition 1) using an encoding tree to model the structure of the semantic space.
Definition 1 (Hierarchical Abstraction).
@@
The hierarchical abstraction of a graph $G$ is formalized by an encoding tree $\mathcal{T}$ that satisfies the following conditions: (1) The root node $\lambda$ in layer 0 contains all nodes in $G$, $\mathcal{T}_{\lambda}=V$. Each node $\alpha\in\mathcal{T}$ represents a partition of responses $\mathcal{T}_{\alpha}\subseteq V$. For any leaf node $\gamma$, $\mathcal{T}_{\gamma}$ contains a single response from $V$. (2) Each non-leaf node $\alpha$ has a nonempty set of immediate successors denoted as $\beta_{0},\beta_{1},\ldots,\beta_{l}$. The collection of subsets $\{\mathcal{T}_{\beta_{0}},\mathcal{T}_{\beta_{1}},\ldots,\mathcal{T}_{\beta_{l}}\}$ forms a partition of $\mathcal{T}_{\alpha}$.
[Li and Pan, 2016] proposed structural entropy for measuring structural information embedded in graph data. Mathematically, the one-dimensional structural entropy of graph $G$ equals to the Shannon entropy of the stationary distribution induced by vertex degrees. It is defined as follows:
@@
$H^{\mathcal{T}}(G;\alpha)=-\frac{g_{\alpha}}{\text{vol}(G)}\log_{2}\frac{\mathcal{V}_{\alpha}}{\mathcal{V}_{\alpha^{-}}},$ (5)